# Controlling dynamical systems into unseen target states using machine learning

### v1

@@ Discussion
In this work, we proposed a model-free, data-driven machine learning methodology for smoothly controlling dynamical systems into previously unseen target states, including those with fundamentally different behaviors. By leveraging a parameter-aware machine learning model, we demonstrated the ability to interpolate and extrapolate system dynamics across unexplored parameter regimes. This approach enables transitions to target states that differ significantly from the observed dynamics, offering a novel framework for controlling complex systems. We utilize a stationary prediction of the target dynamic to define a threshold value, such a as the running correlation dimension, to quantify and avoid turbulent behaviors during transitions. By generating multiple non-stationary transition trajectories and filtering for predictions that don’t exhibit turbulence, we ensure smooth transition to the target dynamics. The control approach is generalizable under the condition that the underlying model extrapolates the bifurcation behavior correctly. The next-generation reservoir computing (NGRC) approach used stands out for its exceptional data efficiency, as further confirmed by our experiments on the Lorenz system. Compared to traditional reservoir computing, NGRC requires more than an order of magnitude less training data while maintaining accurate predictions[30]. This makes it particularly suitable for scenarios where data is sparse or expensive to collect. Additionally, NGRC’s concise architecture allows for hardware implementation on platforms such as FPGAs[35]. The current state in this field allows for NGRC hardware implementations with a single control target state. Notably, the method presented here extends the number of achievable target states per hardware implementation and enables non-stationary control, a key advantage for edge computing applications where real-time adaptability is crucial. However, several limitations warrant further discussion. A notable challenge of the underlying NGRC model is its susceptibility to stability issues during prediction if the hyperparameters are not well optimized. This sensitivity can pose challenges to its adaptability in more complex scenarios. Future work will focus on this issue by refining the two core building blocks of reservoir computing, the reservoir itself and its optimization algorithm, to further progress its adaptability while maintaining high data efficiency with architectural simplicity. Additionally, exploring the integration of NGRC with other machine learning paradigms and testing its performance across a wider range of more complex dynamical systems will further validate its applicability and uncover new potential use cases. By addressing these challenges, the proposed methodology holds promise as a transformative tool for controlling complex dynamical systems, opening pathways to applications where traditional methods fall short.
@@ Methods
Next-generation reservoir computing (NGRC) is a machine learning framework that enhances traditional reservoir computing by leveraging deterministic structures derived from nonlinear vector autoregression (NVAR). Instead of utilizing a randomly weighted internal network, NGRC constructs its feature vectors through a library of unique monomials of time-shifted input variables from the input data. This process involves concatenating $k$ past time series data points, each separated by $s$ time steps, to form a linear time-delay state vector. The final feature vector $\mathbf{r}_{i}$ is created by generating unique monomials of orders $O$ from this state vector, introducing nonlinearity into the model.[25] This way given a d-dimensional time series $\mathbf{X}$, a matrix of feature vectors $\mathbf{R}(\mathbf{X},k,s,O)=(\mathbf{r}_{1},\mathbf{r}_{2},\ldots,\mathbf{r}_{n})$ can be generated. Training is simplified by optimizing only the output layer $\mathbf{W}_{out}$ which maps the feature vectors onto the difference between the current point and the next point in time using ridge regression, such that
@@
To parameterize the feature vector of the NGRC, the approach introduced in [29] is followed, wherein the parameter $\theta$ multiplied by a scaling factor $\gamma$ is added to each element of the feature vector, resulting in $\mathbf{r}_{i}^{\prime}=\mathbf{r}_{i}+\gamma\theta$. Additionally, $\mathbf{r}_{i}^{\prime}$ is expanded by including its own powers up to order $O_{\text{states}}$ to generate the parametrized feature vector $\widetilde{\mathbf{r}}_{i}$. To employ this architecture for multiple stationary dynamics during training, the concept of multifunctionality is leveraged [36]. Therefore, for each time series $\mathbf{X}_{m}$ and its parameter $\theta_{m}$, the parametrized feature vectors $\widetilde{\mathbf{R}}_{m}=(\widetilde{\mathbf{r}}_{m,1},\widetilde{\mathbf{r}}_{m,2},\ldots,\widetilde{\mathbf{r}}_{m,n})$ are generated along with their corresponding training targets $\mathbf{Y}_{m}$. To optimize the output layer with this setup, the parametrized feature matrices $\widetilde{\mathbf{R}}_{m}$ and the training target matrices $\mathbf{Y}_{m}$ are concatenated into single matrices $\widetilde{\mathbf{R}}_{M}$ and $\mathbf{Y}_{M}$, respectively. This allows the output layer $\mathbf{W}_{\text{out}}$ to be learned via ridge regression by solving the optimization
@@
In this way, both stationary and non-stationary dynamics can be predicted, even for parameter ranges that were not seen during training. The results were obtained using a NGRC architecture with parameters $k=2$, $s=47$, $O=[1,2,3]$, $O_{\text{states}}=3$, and $\gamma=0.18$. Additionally a bias feature of $1$ was included in the parameterized feature vector.
@@
Control theory refers to the study and application of methods for influencing the behavior of dynamical systems such that a desired target state is achieved. In the context of chaotic systems, these approaches are able to stabilize unstable periodic orbits by applying small perturbations through external forces. Traditional methods, such as OGY control [37] and delayed feedback control [38], typically rely on phase space techniques, requiring detailed knowledge of the system’s underlying equations or extensive datasets. These approaches are typically limited to steering systems toward simple dynamical target states. While various extensions have been proposed [39], including methods to "chaotify" periodic or synchronized dynamics [40], the ability to reach arbitrary dynamical target states has only recently been demonstrated [28]. In further research [27], this novel approach was used to compare traditional reservoir computing with next-generation reservoir computing (NGRC) in terms of the data requirements for effective control. NGRC demonstrated superior performance, requiring ten times less training data while achieving comparable control capabilities.
@@
In this paper, we build on the control mechanism introduced in [28, 27] to guide dynamical systems into previously unseen target states. This general framework is not restricted to the next-generation reservoir computer employed here and can also be implemented with alternative methods, such as the minimal reservoir computer, traditional reservoir computing techniques, or other suitable machine learning techniques. The core idea is to train the machine learning method to predict the desired dynamics of the system, represented as the state $\textbf{X}(\rho_{i})$ at a given system parameter $\rho_{i}$. Notably, the dynamics of the desired state $\textbf{X}(\rho_{i})$ do not need to be recorded beforehand for training. Instead, the system can leverage the ability to predict unseen states, as explained in the previous Section, and be trained using the available data. If the system instead evolves to a different state Y — due to external influences or changes in its bifurcation parameters $\rho_{j}$ — an external force $\textbf{F}(t)$ is applied to guide the dynamics back to the desired state $\textbf{X}(\rho_{i})$. Without this intervention, the system would persist in state Y, following its natural trajectory $\textbf{u}(t)$.
@@
To compute the appropriate control force $\textbf{F}(t)$, it is necessary to predict how the system trajectory $\textbf{u}(t)$ would evolve if it were in the desired state $\textbf{X}(\rho_{i})$. This hypothetical trajectory is defined as $\textbf{v}(t)$. Assuming the reservoir computer has been trained effectively, it can accurately forecast the future trajectory $\textbf{v}(t)$ as if the system were in state $\textbf{X}(\rho_{i})$. By comparing this predicted trajectory to the actual system evolution in state Y, the control force is computed based on the difference between the actual and hypothetical trajectories:
@@
where $K$ is a scaling parameter for the force which is set to $K=200$. This control force depends only on the observed actual system coordinates and the predicted coordinates, requiring no knowledge of the system’s underlying equations. This feature makes the approach highly versatile and applicable to real-world problems where precise mathematical models are unavailable. For a thorough analysis, we demonstrate this control strategy on a mathematical example system simulated using its known equations instead of observing a real system. In this case, the control force is incorporated by adding it to the system’s differential equations $\dot{f}$. The updated equations are then solved at each time step, incorporating the control force into the system’s evolution
@@
To ensure a smooth transition to the system’s target state without turbulence occurring, a prediction of the target state must be selected that does not exhibit turbulent behavior. Multiple prediction realizations are generated for a given target state, and their correlation dimension is calculated. If the correlation dimension of a prediction is below the $rcd$-cutoff value, which represents the highest correlation dimension of the stationary target dynamics, it is considered free of turbulence and can be used for control. Depending on the reservoir computing method used, various strategies can generate multiple prediction realizations for a given target state. In our approach, we varied the initial conditions slightly and selected the turbulence-free prediction that starts closest to the system’s current coordinates when applying the control force.
Lorenz System
We selected the Lorenz system [41] as an example to demonstrate the functionality of the control mechanisms. In the analyzed bifurcation parameter range of $\rho$, the Lorenz system exhibits periodic behavior for $\rho>100$, undergoes period doubling for $\rho<100$ and transitions to chaotic behavior. The system of equations $\dot{f}(\mathbf{x}(t,\rho))$ is given by
$\dot{x}=\sigma(y-x),\\ \dot{y}=x(\rho-z)-y,\\ \dot{z}=xy-\beta z,$ (6)
where the parameters are set to $\sigma=10$ and $\beta=\frac{8}{3}$. These equations are solved for the simulation results using the fourth-order Runge-Kutta method with a time step size of $\Delta t=0.005$.
@@
To determine whether the system undergoes a smooth transition during the non-stationary process, we employ a running correlation dimension approach computed over a defined window length. This technique captures the system’s dynamic spatial complexity over time. For periodic dynamics, the running correlation dimension values remain close to 1, whereas they exceed 1 when the dynamics exhibit chaotic behavior. In the case of the chaotic target dynamics of the example Lorenz system with $\rho=99.5$, this value remains below $1.21$. During non-stationary simulation, prediction, and control, we detect a non-smooth transition if the running correlation dimension surpasses $1.21$, indicating possible turbulence, while a smooth transition is considered if it remains below this threshold. For each windowed trajectory the correlation dimension $\nu$ quantifies the fractal dimensionality of the trajectory’s phase space [42].
@@
where $\theta$ is the Heaviside function. The states are considered to be close to each other if their distance is less than the threshold distance $r$. The dimension $\nu$ is determined from the scaling relation:
@@
Appendix
Acknowledgements
D. K. gratefully acknowledge the funding provided by Allianz Global Investors (AGI).

### v2

@@ Discussion
In this work, we introduce a model-free and data-driven machine learning methodology to control dynamical systems in previously unseen target states, with a prediction evaluation and selection scheme enabling complex control strategies. Using parameter-aware next-generation reservoir computing, we demonstrate the ability to utilize its extrapolation capabilities to efficiently generate control scenarios that are selected when a predefined control objective is met. We show that this way transient behavior can be suppressed across transition times in a non-stationary control application. Furthermore, we demonstrate that the method can navigate transitions even in scenarios where system collapse is observed frequently, while ensuring fast transitions and avoiding prolonged transient behavior. These scenarios are validated for additive noise on the dynamics of the power system. NGRCs stand out for their exceptional data efficiency. Our experiments confirm this by requiring more than an order of magnitude less training data than results reported for learning the Lorenz system using traditional reservoir computers [25]. This makes the approach particularly suitable for scenarios where data is scarce or expensive to collect. Despite this efficiency, NGRCs require careful hyperparameter optimization to avoid stability issues during prediction. This sensitivity is greater than in other reservoir computing approaches mainly because of these following three factors in combination: the lack of inherent boundedness can cause instability, the feature vector dimension scales combinatorially with time delays and monomial order, and higher dimensionality often fails to improve performance. Once a suitable configuration is achieved, however, the parameter-aware NGRC can generate stable non-stationary predictions and be used as a controller over transients to unseen parameter regions. Recent work, demonstrated that NGRC-based control of dynamical systems is feasible on platforms such as FPGAs [23]. Our method extends this work through parameter awareness, enabling complex control strategies and arbitrary target states beyond the learned parameter range while maintaining equivalent computational requirements within a single trained NGRC architecture. This efficiency, combined with low training data requirements, makes the approach especially suited for edge applications where real-time adaptability is key. Future work will focus on the sensitivity issue by refining the reservoir and its optimization algorithm to further progress its adaptability while maintaining high data efficiency with architectural simplicity. By addressing these challenges, the proposed methodology shows promise as an innovative tool for controlling dynamical systems in complex scenarios and edge environments, opening pathways to applications where traditional methods fall short.
@@ Methods
Next-generation reservoir computing (NGRC) is a machine learning framework that enhances traditional reservoir computing by leveraging deterministic structures derived from non-linear vector autoregression (NVAR). Instead of utilizing a randomly weighted internal network, NGRC constructs its feature vectors through a library of unique monomials of time-shifted input variables from the input data. This process involves concatenating $k$ past time series data points, each separated by $s$ time steps, to form a linear time-delay state vector. The final feature vector $\mathbf{r}_{i}$ is created by generating unique monomials of orders $O$ from this state vector, introducing non-linearity into the model [14]. In this way, given a time series of dimensions d $\mathbf{X}$, a matrix of feature vectors $\mathbf{R}(\mathbf{X},k,s,O)=(\mathbf{r}_{1},\mathbf{r}_{2},\ldots,\mathbf{r}_{n})$ can be generated. Training is simplified by optimizing only the output layer $\mathbf{W}_{out}$ that maps the feature vectors onto the difference between the current point and the next point in time using ridge regression, such that
@@
To parameterize the NGRC feature vector, the approach introduced in [27] is followed, where the parameter $\theta$ multiplied by a scaling factor $\gamma$ is added to each element of the feature vector, resulting in $\mathbf{r}_{i}^{\prime}=\mathbf{r}_{i}+\gamma\theta$. Furthermore, $\mathbf{r}_{i}^{\prime}$ is expanded by including its own powers up to order $O_{\text{states}}$ to generate the parameterized feature vector $\widetilde{\mathbf{r}}_{i}$. To employ this architecture for multiple stationary dynamics during training, the concept of multifunctionality is leveraged [13]. Therefore, for each time series $\mathbf{X}_{m}$ and its parameter $\theta_{m}$, the parametrized feature vectors $\widetilde{\mathbf{R}}_{m}=(\widetilde{\mathbf{r}}_{m,1},\widetilde{\mathbf{r}}_{m,2},\ldots,\widetilde{\mathbf{r}}_{m,n})$ are generated along with their corresponding training targets $\mathbf{Y}_{m}$. To optimize the output layer with this setup, the parameterized feature matrices $\widetilde{\mathbf{R}}_{m}$ and the training target matrices $\mathbf{Y}_{m}$ are concatenated into single matrices $\widetilde{\mathbf{R}}_{M}$ and $\mathbf{Y}_{M}$, respectively. This allows the output layer $\mathbf{W}_{\text{out}}$ to be learned via ridge regression by solving the optimization
@@
In this way, both stationary and non-stationary dynamics can be predicted, even for parameter ranges that were not seen during training.
@@
Control theory refers to the study and application of methods for influencing the behavior of dynamical systems such that a desired target state is achieved. In the context of chaotic systems, these approaches are able to stabilize unstable periodic orbits by applying small perturbations through external forces. Traditional methods, such as OGY control [38] and delayed feedback control [39], typically rely on phase space techniques, which require detailed knowledge of the system’s underlying equations or extensive datasets. These approaches are typically limited to steering systems toward simple dynamical target states. Although various extensions have been proposed [4], including methods to "chaotify" periodic or synchronized dynamics [42], the ability to reach arbitrary dynamical target states has only recently been demonstrated [18]. In further research [17], this novel approach was used to compare traditional reservoir computing with next-generation reservoir computing (NGRC) in terms of data requirements for effective control. NGRC demonstrated superior performance, requiring ten times less training data while achieving comparable control capabilities.
@@
In this paper, we build on the control mechanism introduced in [18, 17] to guide dynamical systems into previously unseen target states. This general framework is not restricted to the next-generation reservoir computer employed here and can also be implemented with alternative methods, such as the minimal reservoir computer, traditional reservoir computing techniques, or other suitable machine learning techniques. The core idea is to train the machine learning method to predict the desired dynamics of the system, represented as the state $\textbf{X}(\rho_{i})$ at a given system parameter $\rho_{i}$. In particular, the dynamics of the desired state $\textbf{X}(\rho_{i})$ does not need to be recorded prior to training. Instead, the system can leverage the ability to predict unseen states, as explained in the previous Section, and be trained using the available data. If the system instead evolves to a different state Y — due to external influences or changes in its bifurcation parameters $\rho_{j}$ — an external force $\textbf{F}(t)$ is applied to guide the dynamics back to the desired state $\textbf{X}(\rho_{i})$. Without this intervention, the system would persist in state Y, following its natural trajectory $\textbf{u}(t)$.
@@
To compute the appropriate control force $\textbf{F}(t)$, it is necessary to predict how the trajectory of the system $\textbf{u}(t)$ would evolve if it were in the desired state $\textbf{X}(\rho_{i})$. This hypothetical trajectory is defined as $\textbf{v}(t)$. Assuming that the reservoir computer has been trained effectively, it can accurately forecast the future trajectory $\textbf{v}(t)$ as if the system were in state $\textbf{X}(\rho_{i})$. By comparing this predicted trajectory with the actual state evolution of the system Y, the control force is calculated based on the difference between the actual and hypothetical trajectories:
@@
where $K$ is a control parameter. This control force depends only on the observed actual system coordinates and the predicted coordinates, which requires no knowledge of the system’s underlying equations. This feature makes the approach highly versatile and applicable to real-world problems where precise mathematical models are unavailable. For a thorough analysis, we demonstrate this control strategy on mathematical example systems simulated using its known equations instead of observing a real system. In this case, the control force is incorporated by adding it to the system’s differential equations $\dot{f}$. The updated equations are then solved at each time step, incorporating the control force into the system’s evolution
$\textbf{u}(t+\Delta{t})=\int_{t}^{t+\Delta{t}}(\dot{f}(\textbf{u}(\tilde{t}))+\textbf{F}(\tilde{t}))d\tilde{t}\$. (5)
To ensure a smooth transition in Lorenz application to the system’s target state without transients occurring, a prediction of the target state must be selected that does not exhibit transient behavior. Multiple prediction realizations are generated for a given target state and their correlation dimension is calculated. If the correlation dimension of a prediction is below the rcd threshold value, where $\text{rcd}<1.21$ which represents the highest correlation dimension of the stationary target dynamics, it is considered free from transients and can be used for control. Depending on the reservoir computing method used, various strategies can generate multiple prediction realizations for a given target state. In this approach, we varied the initial conditions slightly and selected the transient-free prediction that starts closest to the system’s current coordinates when applying the control force. For the power system application with noise, the control method is identical, except for $\mathbf{u(t)}\rightarrow\tilde{\mathbf{u}}(t)=\mathbf{u}(t)+\mathbf{\mu}(t)$ with $\mathbf{\mu}(t)\sim\mathbf{\mathcal{N}}(0,\,\sigma^{2})$. The resulting control force is applied to the dynamical system directly, as well.
Lorenz System
We selected the Lorenz system [33] as an example to demonstrate the functionality of the control mechanisms. In the range of bifurcation parameters analyzed of $\rho$, the Lorenz system exhibits periodic behavior for $\rho>100$, undergoes period doubling for $\rho<100$ and transitions to chaotic behavior. The system of equations $\dot{f}(\mathbf{x}(t,\rho))$ is given by
$\dot{x}=\sigma(y-x),\\ \dot{y}=x(\rho-z)-y,\\ \dot{z}=xy-\beta z,$ (6)
where the parameters are set to $\sigma=10$ and $\beta=\frac{8}{3}$. These equations are solved for the simulation results using the fourth-order Runge-Kutta method with a time-step size of $\Delta t=0.005$. The results were obtained using an NGRC architecture with parameters $k=2$, $s=47$, $O=[1,2,3]$, $O_{\text{states}}=3$ and $\gamma=0.18$. In addition, a bias feature of $1$ was included in the parameterized feature vector. During the control, the control parameter was set to $K=50$ across all transition periods.
Generic power system model
The model is defined by a system of four ordinary differential equations:
$\dot{\delta}_{m}=\omega,$ (7)
$M\dot{\omega}=-d_{m}\omega+P_{m}-E_{m}Y_{m}\sin(\delta_{m}-\delta)V,$ (8)
$K_{qw}\dot{\delta}=-K_{qv2}V^{2}-K_{qv}V+Q(\delta_{m},\delta,V)-Q_{0}-Q_{1},$ (9)
$TK_{qw}K_{pv}\dot{V}=$ $\;K_{pw}K_{qv2}V^{2}+(K_{pw}K_{qv}-K_{qw}K_{pv})V$
$+K_{qw}[P(\delta_{m},\delta,V)-P_{0}-P_{1}]$
$-K_{pw}[Q(\delta_{m},\delta,V)-Q_{0}-Q_{1}]$ (10)
The real power demand $P$ and reactive power demand $Q$ are defined as:
$P(\delta_{m},\delta,V)$ $=-E^{\prime}_{0}Y^{\prime}_{0}V\sin\delta+E_{m}Y_{m}V\sin(\delta_{m}-\delta)$ $\qquad Q(\delta_{m},\delta,V)$ $=-E^{\prime}_{0}Y^{\prime}_{0}V\cos\delta-(Y^{\prime}_{0}+Y_{m})V^{2}+E_{m}Y_{m}V\cos(\delta_{m}-\delta)$
In this formulation, the real‑power demand $P$ and reactive‑power demand $Q$ drive the differential equations governing the load voltage $V$ and the motor angle $\delta$. The term $\delta_{m}$ denotes the angle separation of the rotor between the two generators, while $\omega$ represents the speed of the generator rotor. For a comprehensive technical discussion, see Dobson et al. [12]. The parameters are selected according to the conventions used in previous reservoir computing models [28, 27].
$E^{\prime}_{0}$ $\,=\frac{E_{0}}{\sqrt{1+C^{2}Y_{0}^{-2}-2CY_{0}^{-1}\cos\theta_{0}}}$ $\qquad Y^{\prime}_{0}$ $\,=Y_{0}\sqrt{1+C^{2}Y_{0}^{-2}-2CY_{0}^{-1}\cos\theta_{0}}$ $\qquad\theta^{\prime}_{0}$ $\,=\theta_{0}+\tan^{-1}\!\bigl(\tfrac{CY_{0}^{-1}\sin\theta_{0}}{\,1-CY_{0}^{-1}\cos\theta_{0}\,}\bigr)$ (11)
The parameter values are:
@@
[…]
The parameter $Q_{1}$ acts as the bifurcation parameter, representing the reactive power demand of the system. The bifurcation diagram is produced by integrating the governing equations with a fourth‑order Runge–Kutta (RK4) scheme. Each simulation begins from the initial state
$\mathbf{x}_{0}=(\delta_{m,0},\omega_{0},\delta_{0},V_{0})^{\mathsf{T}}=(0.17,\,0.05,\,0.05,\,0.83)^{\mathsf{T}},$
and is advanced for 20 00020\,000 time steps with a fixed step size of $\Delta t=0.05$. The bifurcation parameter is incremented by $\Delta Q_{1}=2\times 10^{-6}$ in the analysis of the bifurcation diagrams. The training data are obtained for the following values of the bifurcation parameter:
$Q_{1}\in\{2.98947,\,2.98952,\,2.98957,\,2.98962,\,2.98967,\,2.98972,\,2.98977\}.$
The hyperparameters of the trained NGRC are $k=2$, $s=63$, $O=[1,2]$, $O_{\text{states}}=2$ with an additional bias term, and $\gamma=0.292$. For the non-stationary predictions of the instantaneous parameter change, we hold $Q^{initial}_{1}$ for 250 seconds constant and then switch to $Q^{target}_{1}$ for additional 500 seconds.
Prediction evaluation and selection scheme
Control scenarios are predicted with Eq. 3. The vector $\boldsymbol{\theta}$ specifies the desired non‑stationary dynamics, and its length sets the length of the prediction of the NGRC. For each element in $\mathbf{\theta}$, the NGRC predicts the next point in time under the current $\theta_{i}$. Varying the NGRC warm‑up phase yields distinct non‑stationary trajectories. Each trajectory is evaluated against a predefined objective- for example, the running correlation dimension of the trajectory must remain below a threshold or exhibiting short transition periods measured by the spectral entropy over time. A trajectory that meets the objective is adopted as the control strategy. In the applications discussed, we start from $\theta_{initial}$ for some time to control the system from its initial state to the initial state of the prediction before the parameter change is forced.
@@
To determine whether the system undergoes a smooth transition during the non-stationary process, we employ a running correlation dimension approach computed over a defined window length. This technique captures the dynamic spatial complexity of the system over time. For periodic dynamics, the running correlation dimension values remain close to 1, whereas they exceed 1 when the dynamics exhibit chaotic behavior. In the case of the chaotic target dynamics of the example Lorenz system with $\rho=99.5$, this value remains below $1.21$. During non-stationary simulation, prediction, and control, we detect a non-smooth transition if the running correlation dimension surpasses $1.21$, indicating possible transients, while a smooth transition is considered if it remains below this threshold. For each windowed trajectory, the correlation dimension $\nu$ quantifies the fractal dimensionality of the trajectory’s phase space [15]
@@
where $\theta$ is the Heaviside function. The states are considered close to each other if their distance is less than the threshold distance $r$. The dimension $\nu$ is determined from the scaling relation
@@
Spectral Entropy Measure
The spectral entropy metric is used to detect intermittent episodes in the non‑stationary power‑system model, capturing the transition from one periodic regime to another. We evaluate the short-time Fourier transform [47] (STFT) calculated with a window of $n_{\text{perseg}}=2000$ samples and an overlap of $n_{\text{overlap}}=1950$ samples, i.e. 97. 5%. Only components in the $0.1\text{–}2\penalty 10000\ \text{Hz}$ band are retained for subsequent processing, as this range is known to carry the dynamics of interest in the system analyzed. Frames whose center times violate $n_{\text{perseg}}$ at the start and end of the analyzed time series are discarded. For every remaining STFT frame $\ell$ the band‑limited power spectrum $P_{k\ell}=|Z_{xx}(f_{k},t_{\ell})|^{2}$ is normalized to the unit mass and its Shannon entropy[44] is evaluated,
$H_{\ell}\;=\;-\sum_{k}P_{k\ell}\,\ln P_{k\ell},$ (14)
yields the trace of entropy $H_{\ell}$ against the center time $t_{\ell}$. A first-order difference $\Delta H_{\ell}=H_{\ell}-H_{\ell-1}$ is used as a discrete approximation of the temporal derivative. The onset of an intermittent episode is flagged at the earliest time index for which $\Delta H_{\ell}$ exceeds a fixed rise threshold $\Delta H_{\text{rise}}=+0.01$. The episode is considered to end with the first subsequent index where $\Delta H_{\ell}$ falls below the corresponding fall threshold $\Delta H_{\text{fall}}=-0.01$. The duration of the detected intermittent period is adjusted to match the number of segments used in the analysis.
@@
Data and Code Availability
The data and code are available from the corresponding author upon reasonable request.
Acknowledgments
D. K. gratefully acknowledges the funding provided by Allianz Global Investors.
Declaration of Interests
The German Aerospace Center, together with A.H., have filed a patent application related to this work.